# Stealthy Patch-Wise Backdoor Attack in 3D Point Cloud via Curvature Awareness

### v1

@@ 3.3 Patch-wise Spectral Trigger Injection
To overcome these limitations, we propose designing local patch-wise triggers in the spectral domain [12, 20]. Specially, a selected point cloud patch $S^{\prime}\in\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ is first transformed into an unweighted KNN graph $G=(V,A)$. Here $V\in\mathbb{R}^{k_{p}\times 3}$ represents the set of vertices, where each vertex corresponds to a point in the patch $S^{\prime}$ and connects to its $k_{g}$-nearest neighbors. The adjacency matrix $A\in\mathbb{R}^{k_{p}\times k_{p}}$ encodes connectivity among points, where $A_{ij}=1$ if point $i$ and $j$ are connected otherwise $A_{ij}=0$. From this, we compute the combinatorial graph Laplacian matrix [29] defined as $L=D-A$, where $D\in\mathbb{R}^{k_{p}\times k_{p}}$ is the degree matrix, with each entry defined as $D_{ij}=\sum_{j=1}^{N}A_{ij}$. Since $L$ is real, symmetric, and positive semi-definite, we perform eigen decomposition on it as $L=U\Lambda U^{T}$, where $U\in\mathbb{R}^{k_{p}\times k_{p}}$ is an orthonormal matrix containing the eigenvectors and $\Lambda$ consists of eigenvalues.
@@
where $\hat{S}\in\mathbb{R}^{k_{p}\times 3}$. Correspondingly, the Inverse GFT (IGFT) could reconstruct the spatial feature from the spectral domain:
@@
Figure 3: Comparison of original and perturbated patches. The second row shows the spectral features of the original patch and spectral perturbations across different spectrum bands.
As illustrated in Figure 3, low-spectrum perturbations cause significant deformation to the patch of empennage, while high-spectrum perturbations introduce excessive outliers. Perturbations applied within intermediate spectral bands mitigate outliers while preserving the overall shape, presenting a key challenge for optimal spectral trigger design. To address this, we propose an optimizable spectral trigger $\xi\in\mathbb{R}^{k_{p}\times 3}$, which is applied to the spectral features of selected patches to generate the poisoned patches $S^{p}$:
@@
where $S^{\prime}\in\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$. The final poisoned point cloud sample $P^{\prime}$ is then constructed by replacing the original selected patches with the reconstructed patches while keeping the unselected regions unchanged:
$P^{\prime}=\mathcal{T}(P)=(P\setminus\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\})\cup\{S_{1}^{p},S_{2}^{p},...,S_{m}^{p}\}.$ (7)
Figure 4: Visual comparison of poisoned samples from different backdoor attack methods. SPBA preserves structural integrity and ensures smooth regions (e.g., airplane body, vase) remain undisturbed, enhancing imperceptibility.
@@ 3.4 Optimization Strategy
$\begin{split}\!\!\!\!L_{M}=\!\!\!\!\sum_{(P,Y)\in D_{b}}\!\!\!\!\!\!L_{CE}&(f_{\theta}(P),Y)+\!\!\!\!\!\!\!\!\sum_{(P^{\prime},Y_{t})\in D_{p}}\!\!\!\!\!\!\!\!L_{CE}(f_{\theta}(P^{\prime}),Y_{t}),\!\!\!\!\!\!\end{split}$ (8)
@@
$L_{T}=\sum_{(P,Y)\in D}\left(L_{CE}(f_{\theta}(\mathcal{T}(P)),Y_{t})+L_{reg}(P,\mathcal{T}(P))\right).$ (9)
In the regularizer function $L_{reg}$, we employ Euclidean distance (ED) and Chamfer distance (CD) [36] to enforce similarity between the original and poisoned point clouds, and Hausdorff distance (HD) [33] is used to control the worst scenario of outliers. The final regularization term is formulated as:
$\begin{split}L_{reg}=\lambda_{1}L_{ED}(P,\mathcal{T}(P))+\lambda_{2}L_{CD}(P,\mathcal{T}(P))\\ +\lambda_{3}L_{HD}(P,\mathcal{T}(P)),\end{split}$ (10)
@@ 4.1 Experiment Details
Datasets and Models. We evaluate our proposed method on two widely used 3D point cloud classification datasets: 1) ModelNet40 [34]: a benchmark dataset containing 40 object categories, with an official split of 9,843 training samples and 2,468 testing samples. 2) ShapeNetPart [40]: a subset of ShapeNet, consisting of 12,128 training samples and 2,874 testing samples across 16 object categories. For both datasets, we uniformly sample 1,024 points from each shape along with their normal vectors, normalizing the coordinates for consistency. For backdoored models, we adopted the most widely used models in previous work [13, 7], including PointNet [24], PointNet++ [25], DGCNN [31], and PointNext [26].
Attacks Setup. We compare our proposed SPBA method against four representative point cloud attack methods: PointBA-I [13], PointBA-O [13], IRBA [7], and IBAPC [5]. To reproduce PointBA-I, we implant a spherical cluster comprising 3% of the total points, with a fixed radius of 0.05, positioned at $(0.5,0.5,0.5)$ in the coordinate space. For PointBA-O, we apply a $10^{\circ}$ rotation along the Z-axis. For IRBA and IBAPC, we follow the default configurations specified in their original papers. For our proposed SPBA, $g$ and $k_{p}$ are both set as 32 for patch decomposition, while the values of $k_{c}$ and $k_{g}$ are both set as 10 for the KNN graph construction. And $m=16$ patches are selected for trigger injection. For the regularizer function in Eq. 10, we set $\lambda_{1}=1$, $\lambda_{2}=5$ and $\lambda_{3}=1$. However, for PointNet, we adjust $\lambda_{1}$ to 0.1 to improve stability. For all attack methods, we set the poisoning rate $\rho=0.1$. The target class is chosen as Chair ($Y_{t}=8$) for ModelNe40 and Lamp ($Y_{t}=8$) for ShapeNetPart. Two Adam optimizers are employed for backdoored models traning and spectral trigger optimization, with learning rates of 0.001 and 0.01, respectively.
@@
Table 1: Comparison of attack performance across different models on ModelNet40. Both BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000. The best results are bolded and the second-best results are underlined.
Model Clean PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
PointNet 89.14 89.67 99.03 4.08 88.17 94.77 5.31 88.25 93.27 18.54 88.42 94.15 3.58 88.90 97.89 2.60
PointNet++ 90.80 90.36 98.87 4.08 89.55 96.68 5.31 90.40 98.26 18.54 90.72 96.11 0.47 89.42 97.97 0.11
DGCNN 89.75 91.05 100.00 4.08 89.67 95.02 5.31 89.55 92.38 18.54 88.04 97.44 0.54 89.75 97.61 0.69
PointNext 89.22 88.21 95.83 4.08 88.94 96.31 5.31 88.41 92.46 18.54 88.70 97.69 1.09 88.62 96.85 0.92
Average 89.73 89.82 98.43 4.08 89.08 95.69 5.31 89.15 94.09 18.54 88.97 96.35 1.42 89.17 97.58 1.08
Table 2: Comparison of attack performance across different models on ShapeNetPart. Both BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000. The best results are bolded and the second-best results are underlined.
Model Clean PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
PointNet 98.47 98.30 99.72 7.48 97.95 98.34 5.67 98.12 96.80 16.06 97.98 95.86 1.82 98.30 98.40 1.22
PointNet++ 98.99 98.75 99.69 7.48 98.50 98.78 5.67 98.89 99.03 16.06 98.47 98.92 0.08 98.26 98.30 0.02
DGCNN 99.06 98.92 100.00 7.48 98.85 98.50 5.67 98.68 97.63 16.06 97.36 98.68 0.75 98.43 99.23 0.44
PointNext 98.82 98.61 99.51 7.48 98.10 96.83 5.67 96.31 96.17 16.06 97.32 98.37 0.67 98.53 98.45 0.37
Average 98.83 98.64 99.73 7.48 98.35 98.11 5.67 98.00 97.41 16.06 97.78 97.96 0.83 98.38 98.60 0.51
@@ 4.2 Attack Effectiveness
As shown in Table 1, our proposed SPBA achieves an average BA of 89.02%, which is comparable to the clean model and exceeds most attack methods, ensuring that the model retains its normal classification performance. Although PointBA-I achieves the highest average ASR of 98.43%, it comes at the cost of a significantly higher average CD of 4.08, making it much more detectable. In contrast, SPBA maintains the lowest average CD of 1.08, highlighting its superior imperceptibility. What’s more, SPBA achieves an average ASR of 97.58%, surpassing transform-based attacks PointBA-O and IRBA by 1.89% and 3.49% respectively. Compared to the global spectral attack IBAPC, SPBA not only achieves higher BA and ASR on average but also further reduces average CD by 0.34, demonstrating the stealthiness of injecting an optimized spectral trigger into geometrically complex patches. Table 2 shows more experiment results on ShapeNetpart. SPBA achieves the second-highest average ASR and BA while maintaining the lowest average CD, achieving the best balance between attack effectiveness and stealthiness. These results further establish SPBA as a more effective and imperceptible backdoor attack method compared to existing approaches.
Table 3: Resistance of backdoor attacks against varied data augmentations
@@
No 90.36 98.87 89.55 96.68 90.40 98.26 90.72 96.11 89.42 97.97
Rotation 89.42 98.95 86.83 7.66 90.52 97.69 89.91 95.95 89.34 98.22
Rotation3D 89.71 99.11 88.78 30.51 90.68 92.79 90.24 95.34 89.38 97.24
Scaling 90.72 99.55 91.65 96.92 91.17 99.47 91.49 97.08 91.57 98.26
Shift 90.11 99.03 89.67 96.07 90.03 97.93 90.52 95.99 89.59 97.08
Dropout 90.52 99.39 90.07 96.11 90.40 94.85 91.05 95.02 90.76 97.24
Jitter 89.42 98.70 89.34 95.06 90.07 90.03 90.64 92.18 90.96 95.01
4.3 Attack Stealthiness
We visualize some poisoned samples generated by different backdoor attack methods in Figure 4. As observed, PointBA-I introduces noticeable anomalies compared to clean samples, significantly increasing its detectability. Although PointBA-O and IRBA generate natural-looking poisoned samples by applying transformation, they still cause considerable deformations in shape. Compared to IBAPC, our proposed SPBA maintains the smoothness of regions with simple structures, thereby enhancing the imperceptibility of poisoned samples. For instance, in the flower pot example, SPBA selectively perturbs the flower region while preserving the smooth structure of the vase, making the modifications less perceptible to human observers.
Table 4: Comparison of attack performance between IBAPC and SPBA with varied poisoned patches.
Method IBAPC-L IBAPC-H16 SPBA-H8
[…]
PointNet 61.75 1.93 46.03 1.65 86.10 1.39
PointNet++ 85.66 0.25 66.53 0.22 85.90 0.06
DGCNN 69.00 0.26 46.88 0.28 83.91 0.32
PointNext 57.82 0.61 33.63 0.48 77.47 0.35
Average 68.56 0.76 48.27 0.66 83.35 0.53
4.4 Comparison of Spectral Triggers
Both SPBA and IBAPC involve spectral trigger design, but differ in their scope. SPBA applies a patch-wise spectral trigger independently to the 16 patches with the highest PSI, whereas IBAPC employs a global spectral trigger that affects the entire sample. To analyze the impact of perturbation localization, we evaluate the well-trained spectral trigger on the ModelNet40 test dataset using PointNet++, which only affects the ASR and CD of backdoored models on the poisoned data. Specially, we decompose the trigger of IBAPC into IBAPC-H16 (confined to complex regions) and IBAPC-L (confined to smooth regions). As shown in Table 4, IBAPC-L achieves a 20.29% higher average ASR than IBAPC-H16 under near-intensity perturbations. This result suggests that IBAPC relies more on smooth-region perturbations, which reduce imperceptibility, as illustrated in Figure 4. For a fair comparison, we introduce SPBA-H8, a weaker test version of our method, which applies the patch-wise trigger to only eight complex patches during inference, further minimizing perturbation intensity. However, SPBA-H8 still surpasses IBAPC-L and IBAPC-H16 in ASR by 14.79% and 35.08%, respectively, while maintaining the lowest CD. This highlights the effectiveness of patch-wise spectral triggers in achieving an optimal balance between attack performance and stealthiness.
@@ 4.5 Resistance to Defense Methods
Resistance to Data Augmentation. Since data augmentation is widely employed to enhance the robustness of classification models, we introduce six types of augmentations to evaluate their impact on backdoor attack performance: 1) Rotation: random rotation around the z-axis with a maximum angle of $10^{\circ}$; 2) Rotation3D: random rotation up to $10^{\circ}$ along all three axes; 3) Scaling: scaling within a range of 0.5 to 1.5; 4) Shift: shifting randomly between -0.1 and 0.1 along each axis; 5) Dropout: removing up to 20% of points; and 6) Jitter: introducing noise sampled from $N(0,0.02)$. The experiments are conducted on the ModelNet40 dataset using the PointNet++ model. As shown in Table 3, different augmentations affect attack methods to varying degrees. Due to the high visibility of the trigger of PointBA-I, augmentations have minimal impact, as the trigger remains distinct. However, for rotation-based PointBA-O, applying Rotation significantly reduces attack effectiveness. For IRBA, which relies on local transformations, augmentations such as Rotation3D, Dropout and Jitter cause a notable decline in attack performance, with ASR dropping by up to 8.23%. Conversely, for optimization-based attacks like IBAPC and our proposed SPBA, only Jitter leads to a slight ASR reduction, while other augmentations have negligible effects, demonstrating their robustness against most augmentations.
@@
Resistance to SOR. Statistical Outlier Removal (SOR) [45] is a widely used technique designed to eliminate outliers based on the local neighborhood density, determined by the KNN algorithm. In our implementation, we set KNN to use 40 nearest neighbors and set the standard deviation threshold to 1. To assess the impact of SOR on backdoor attack performance, we apply it during inference on PointNet++ models trained on ModelNet40. Results show that only PointBA-I experiences a significant ASR reduction of 22.94%, as SOR effectively removes its trigger points as outliers. This suggests that SOR serves as an effective defense specifically against injection-based attacks while having limited impact on other attack methods.
Resistance to the Gradient-based Method. As many studies on backdoor attacks assess trigger stealthiness using gradient-based methods [28, 23], we employ a point cloud saliency map [44] to identify salient points that contribute most to the model’s decision-making. Specially, 40 points with the highest saliency scores are marked in red. As illustrated in Figure 5, the spherical trigger of PointBA-I is easily detected. In contrast, our proposed SPBA exhibits a saliency distribution similar to the clean model, indicating enhanced stealthiness and reduced detectability.
@@
Table 5: Comparison of attack performance of SPBA with various patch selection strategies.
Strategy H-PSI L-PSI HD LD Random
[…]
4.6 Ablation Study
Effect of Patch Selection Strategies. In our default setting, we select the 16 patches with the highest patch imperceptibility scores (PIS), denoted as H-PSI. To evaluate the effectiveness of this selection strategy, we compare H-PSI with four alternative patch selection strategies: (1) L-PSI: Select the patches with the lowest PIS. (2) HD: Select the patches with the highest patch densities, computed as the sum of distances from points to their respective patch central points. (3) LD: Select the patches with the lowest patch densities. (4) Random: Randomly select the patches. We evaluate their impact on ModelNet40 using the PointNet++ model. As shown in Table 5, the H-PSI strategy achieves the highest attack effectiveness and the strongest stealthiness. This result underscores the advantage of selecting high-imperceptibility patches for local spectral trigger injection.
5 Conclusion
In this paper, we propose SPBA, a novel backdoor attack method against 3D point cloud models. Unlike existing methods that rely on sample-wise trigger injection, which introduces noticeable geometric distortions, SPBA employs a local patch-wise trigger that selectively perturbs geometrically complex regions. By leveraging Graph Fourier Transform, our proposed patch-wise trigger minimizes geometric deformations while maintaining high attack effectiveness. Experimental results on ModelNet40 and ShapeNetPart demonstrate that SPBA achieves a high attack success rate with minimal perceptibility, outperforming prior sample-wise backdoor attacks in stealthiness. These findings highlight the effectiveness of patch-wise triggers as a promising direction for stealthy and adaptive backdoor attacks in 3D deep learning.
Appendix
Appendix A Ablation Study
A.1 Effect of the Target Class $Y_{t}$
Figure 6: Effect of the target class selection.
A.2 Effect of the Poisoning Rate $\rho$
Figure 7: The effect of the poisoning rate.
Table 6: Effect of selected patch number $m$

### v2

@@ 3.3 Patch-wise Spectral Trigger Injection
To overcome these limitations, we propose designing local patch-wise triggers in the spectral domain [8, 14]. Specially, a selected point cloud patch $S^{\prime}\in\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\}$ is first transformed into an unweighted KNN graph $G=(V,A)$. Here $V\in\mathbb{R}^{k_{g}\times 3}$ represents the set of vertices, where each vertex corresponds to a point in the patch $S^{\prime}$. The adjacency matrix $A\in\mathbb{R}^{k_{g}\times k_{g}}$ is constructed by connecting each point to its $k_{p}$-nearest neighbors, where $k_{p}$ is a hyperparameter for graph construction ($A_{ij}=1$ if point $i$ and $j$ are connected otherwise $A_{ij}=0$).
From this, we compute the combinatorial graph Laplacian matrix [23] defined as L=DAL=D-A, where $D\in\mathbb{R}^{k_{g}\times k_{g}}$ is the degree matrix, with each entry defined as $D_{ij}=\sum_{j=1}^{N}A_{ij}$. Since $L$ is real, symmetric, and positive semi-definite, we perform eigen decomposition on it as $L=U\Lambda U^{T}$, where $U\in\mathbb{R}^{k_{g}\times k_{g}}$ is an orthonormal matrix containing the eigenvectors and $\Lambda$ consists of eigenvalues.
@@
where $\hat{S}\in\mathbb{R}^{k_{g}\times 3}$. Correspondingly, the Inverse GFT (IGFT) could reconstruct the spatial feature from the spectral domain:
@@
Then, we propose an optimizable spectral trigger $\xi\in\mathbb{R}^{k_{g}\times 3}$, which is applied to the spectral features of selected patches to generate the poisoned patches $S^{p}$:
@@
The final poisoned point cloud sample $X^{p}$ is then constructed by replacing the original selected patches with the reconstructed patches while keeping the unselected regions unchanged:
$X^{p}=\mathcal{T}(X)=(X\setminus\{S^{\prime}_{1},S^{\prime}_{2},...,S^{\prime}_{m}\})\cup\{S_{1}^{p},S_{2}^{p},...,S_{m}^{p}\}.$ (7)
@@ 3.4 Optimization Strategy
$L_{M}=\sum_{D_{b}}L_{CE}(f_{\theta}(X),Y)+\sum_{D_{p}}L_{CE}(f_{\theta}(X^{p}),Y_{t}),$ (8)
@@
$L_{T}=\sum_{D}\left(L_{CE}(f_{\theta}(\mathcal{T}(X)),Y_{t})+L_{reg}(X,\mathcal{T}(X))\right).$ (9)
In the regularizer function $L_{reg}$, we employ Euclidean distance $L_{ED}(X,\mathcal{T}(X))$ and Chamfer distance $L_{CD}(X,\mathcal{T}(X))$ to enforce similarity between the original and poisoned point clouds, and Hausdorff distance $L_{HD}(X,\mathcal{T}(X))$ is used to control the worst scenario of outliers. The final regularization term is formulated as:
$L_{reg}=\lambda_{1}L_{ED}+\lambda_{2}L_{CD}+\lambda_{3}L_{HD},$ (10)
@@
Table 1: Comparison of attack performance across different models on ModelNet40. BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000. The best results are bolded and the second-best results are underlined.
Model Clean PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
PointNet 89.14 89.67 99.03 4.08 88.17 94.77 5.31 88.25 93.27 18.54 88.21 91.57 3.04 88.90 97.89 2.60
PointNet++ 90.80 90.36 98.87 4.08 89.55 96.68 5.31 90.40 98.26 18.54 89.67 95.58 0.68 89.42 97.97 0.11
DGCNN 89.75 91.05 100.00 4.08 89.67 95.02 5.31 89.55 92.38 18.54 88.24 95.44 0.87 89.75 97.61 0.69
PointNext 89.22 88.21 95.83 4.08 88.94 96.31 5.31 88.41 92.46 18.54 88.70 96.60 1.48 88.62 96.85 0.92
Average 89.73 89.82 98.43 4.08 89.08 95.69 5.31 89.15 94.09 18.54 88.70 94.80 1.52 89.17 97.58 1.08
Table 2: Comparison of attack performance across different models on ShapeNetPart. BA and ASR are presented as percentages, and CD values in tables are scaled by a factor of 1,000. The best results are bolded and the second-best results are underlined.
Model Clean PointBA-I PointBA-O IRBA IBAPC SPBA
[…]
PointNet 98.47 98.30 99.72 7.48 97.95 98.34 5.67 98.12 96.80 16.06 97.80 96.01 1.73 98.30 98.40 1.22
PointNet++ 98.99 98.75 99.69 7.48 98.50 98.78 5.67 98.89 99.03 16.06 98.43 98.46 0.22 98.26 98.30 0.02
DGCNN 99.06 98.92 100.00 7.48 98.85 98.50 5.67 98.68 97.63 16.06 97.54 95.44 0.80 98.43 99.23 0.44
PointNext 98.82 98.61 99.51 7.48 98.10 96.83 5.67 96.31 96.17 16.06 97.35 95.99 0.44 98.53 98.45 0.37
Average 98.83 98.64 99.73 7.48 98.35 98.11 5.67 98.00 97.41 16.06 97.78 96.48 0.80 98.38 98.60 0.51
@@ 4.1 Experiment Details
Datasets and Models. We evaluate our proposed method on two widely used 3D point cloud classification datasets: 1) ModelNet40 [27]: a benchmark dataset containing 40 object categories, with an official split of 9,843 training samples and 2,468 testing samples; 2) ShapeNetPart [31]: a subset of ShapeNet, consisting of 12,128 training samples and 2,874 testing samples across 16 object categories. For both datasets, we uniformly sample 1,024 points from each shape along with their normal vectors, normalizing the coordinates for consistency. For backdoored models, we adopted the most widely used models in previous work [10, 4], including PointNet [18], PointNet++ [19], DGCNN [25], and PointNext [20].
Attacks Setup. We compare our proposed SPBA method against four representative point cloud attack methods: PointBA-I [10], PointBA-O [10], IRBA [4], and IBAPC [2]. To reproduce PointBA-I, we implant a spherical cluster comprising 3% of the total points, with a fixed radius of 0.05, positioned at $(0.5,0.5,0.5)$ in the coordinate space. For PointBA-O, we apply a $10^{\circ}$ rotation along the Z-axis. For IRBA and IBAPC, we follow the default configurations specified in their original papers. For our proposed SPBA, $g$ and $k_{g}$ are both set as 32 for patch decomposition, while the values of $k_{c}$ and $k_{p}$ are both set as 10 for the KNN graph construction. And $m=16$ patches are selected for trigger injection. For the regularizer function in Eq. (10), we set $\lambda_{1}=1$, $\lambda_{2}=5$ and $\lambda_{3}=1$. However, for PointNet, we adjust $\lambda_{1}$ to 0.1 to improve stability. For all attack methods, we set the poisoning rate $\rho=0.1$. The target class is chosen as Chair ($Y_{t}=8$) for ModelNe40 and Lamp ($Y_{t}=8$) for ShapeNetPart. Two Adam optimizers are employed for backdoored models training and spectral trigger optimization, with learning rates of 0.001 and 0.01, respectively. The training epoch is set to 100. All experiments were conducted on a machine equipped with an NVIDIA RTX 4090 GPU.
@@ 4.2 Attack Effectiveness
The results presented in Tab. 1 provide a comparative evaluation of various backdoor attack methods on ModelNet40. As shown in Tab. 1, our proposed SPBA achieves the lowest average CD value of 1.08 among all attacks, demonstrating its state-of-the-art stealthiness. What’s more, SPBA outperforms most backdoor attacks according to BA and ASR. In particular, compared to the sample-wise spectral attack IBAPC, SPBA achieves notable improvements of 0.47% in average BA and 2.78% in average ASR, highlighting the effectiveness of our patch-wise spectral trigger. Although PointBA-I achieves the highest average ASR of 98.43%, it comes at the cost of a significantly higher average CD of 4.08, making it much more detectable. Subsequent defense experiments further confirm that the trigger used in PointBA-I is especially susceptible to detection or removal by existing defense methods. Tab. 2 shows more experimental results on the ShapeNetpart dataset. SPBA achieves a similar performance, offering the best balance between attack effectiveness and stealthiness. These results further validate the superiority of SPBA as a more effective and imperceptible backdoor attack method compared to existing attacks.
Figure 3: Visual comparison of poisoned samples from different backdoor attack methods. Our proposed SPBA preserves structural integrity and ensures smooth regions (e.g., vase) remain undisturbed, enhancing imperceptibility.
4.3 Attack Stealthiness
We visualize some poisoned samples generated by different backdoor attack methods in Fig. 3. As observed, PointBA-I introduces noticeable anomalies compared to clean samples, significantly increasing its detectability. Although PointBA-O and IRBA generate natural-looking poisoned samples by applying transformation, they still cause considerable deformations in shape. Compared to IBAPC, SPBA selectively perturbs the flower region while preserving the smooth structure of the vase, making the modifications less perceptible to human observers. Additional visualization results are provided in the Supplementary. Furthermore, we conduct a human perceptual study to evaluate the stealthiness of the attack. Following the setting in IRBA [4], a mixture of clean and poisoned samples is presented to 20 participants. Each participant rates the probability of a sample being maliciously modified on a 5-point scale, where lower scores indicate higher stealthiness. Tab. 3 shows that SPBA achieves the best stealthiness with the lowest score.
Table 3: Scores of the human perceptual study.
Method PointBA-I PointBA-O IRBA IBAPC SPBA
Score $\downarrow$ 4.79 3.33 4.05 2.76 1.53
Table 4: Computational cost of spectral attacks.
Method IBAPC FPSP SPBA
MFLOPS $\downarrow$ 939.72 170.06 14.74
4.4 Attack Efficiency
As shown in the Eq. (6), the computational cost of spectral attacks is highly dependent on the trigger size. And we further report the computational cost of different spectral attacks (with a batch size of 64 during the inference stage) in Tab. 4. The prior spectral attack IBAPC uses a large trigger of size $\xi^{\text{IBAPC}}\in\mathbb{R}^{1024\times 3}$, resulting in a substantial computational cost of 939.72 MFLOPs, which severely hinders its efficiency and practicality for deployment. In contrast, although our proposed SPBA perturbs an average of 452 points, which account for 44% of the sample, our patch-based spectral trigger is much smaller ($\xi^{SPBA}\in\mathbb{R}^{32\times 3}$). The computational cost of SPBA is only 14.74 MFLOPs, which is comparable to that of a single fully connected layer. Furthermore, similar to IBAPC, we construct a sample-based spectral trigger using the FPS algorithm (denoted as FPSP, $\in\mathbb{R}^{452\times 3}$) that perturbs the same number of points as SPBA. However, it significantly increases the computational cost by a factor of 11.54 compared to our patch-based trigger. Moreover, it results in weaker attack performance (shown in Tab. 7), highlighting the advantage of our patch-wise design. These results demonstrate the efficiency and effectiveness of our proposed patch-wise spectral trigger.
Table 5: Resistance of backdoor attacks to varied data augmentations.
@@ 4.4 Attack Efficiency
No 90.36 98.87 89.55 96.68 90.40 98.26 89.67 95.58 89.42 97.97
Rotation 89.42 98.95 86.83 7.66 90.52 97.69 90.59 95.90 89.34 98.22
Rotation3D 89.71 99.11 88.78 30.51 90.68 92.79 90.23 95.94 89.38 97.24
Scaling 90.72 99.55 91.65 96.92 91.17 99.47 90.85 96.47 91.57 98.26
Shift 90.11 99.03 89.67 96.07 90.03 97.93 89.74 94.23 89.59 97.08
Dropout 90.52 99.39 90.07 96.11 90.40 94.85 88.37 94.93 90.76 97.24
Jitter 89.42 98.70 89.34 95.06 90.07 90.03 89.91 91.65 90.96 95.01
@@ 4.5 Resistance to Defense Methods
Resistance to Data Augmentation. Since data augmentation is widely employed to enhance the resistance of classification models, we introduce six types of augmentations to evaluate their impact on backdoor attack performance: 1) Rotation: random rotation around the z-axis with a maximum angle of $10^{\circ}$; 2) Rotation3D: random rotation up to $10^{\circ}$ along all three axes; 3) Scaling: scaling within a range of 0.5 to 1.5; 4) Shift: shifting randomly between -0.1 and 0.1 along each axis; 5) Dropout: removing up to 20% of points; and 6) Jitter: introducing noise sampled from $N(0,0.02)$. The experiments are conducted on the ModelNet40 dataset using the PointNet++ model. As shown in Tab. 5, different augmentations affect attack methods to varying degrees. Due to the high visibility of the trigger of PointBA-I, augmentations have minimal impact, as the trigger remains distinct. However, for rotation-based PointBA-O, applying Rotation significantly reduces attack effectiveness. For IRBA, which relies on local transformations, augmentations such as Rotation3D, Dropout and Jitter cause a notable decline in attack performance, with ASR dropping by up to 8.23%. Conversely, for optimization-based attacks like IBAPC and our proposed SPBA, only Jitter leads to a slight ASR reduction, while other augmentations have negligible effects, demonstrating their resistance against most augmentations.
@@
Resistance to SOR. Statistical Outlier Removal (SOR) [36] is designed to eliminate outliers based on the local neighborhood density using the KNN algorithm. In our implementation, we set KNN to use 40 nearest neighbors and set the standard deviation threshold to 1. We apply it during inference on PointNet++ models trained on ModelNet40. The results show that only PointBA-I experiences a significant drop in ASR, decreasing from 98.87% to 75.44%, while the ASR reduction for other attacks remains within 3%. This suggests that SOR serves as an effective defense specifically against injection-based attacks while having a limited impact on other attack methods.
Resistance to PointCRT. PointCRT [9] utilizes a nonlinear classifier to detect the abnormal corruption resistance of poisoned samples. The F1 score (F1) and Area Under the Curve (AUC) are employed to evaluate the ability of PointCRT to detect the attack, where lower values indicate that the attack is more resistant to detection by PointCRT. As shown in Tab. 6, SPBA outperforms almost all the attacks except PointBA-O, which exhibits limited attack performance and is vulnerable to data augmentations.
Table 6: Detect accuracy of PoinCRT on varies attacks.
Method PointBA-I PointBA-O IRBA IBAPC SPBA
F1 $\downarrow$ 95.94 72.89 93.98 92.84 87.76
AUC $\downarrow$ 97.52 77.76 95.75 92.93 87.89
@@
Table 7: Comparison of attack performance of SPBA with various selection strategies.
Method L-PSI Random FPSP H-PSI(default)
BA $\uparrow$ 87.99 88.19 87.84 89.17
ASR $\uparrow$ 95.53 96.35 96.45 97.58
CD $\downarrow$ 1.16 1.29 1.39 1.08
Table 8: Effect of selected patch number $m$.